OpenStudy (lfghjyt):

)How is the expression (5 + y) ⋅8 described by its factors and terms? It has two factors and the factor (5 + y) has one term. It has three factors and the factor (5 + y) has one term. It has three factors and the factor (5 + y) has two terms. It has two factors and the factor (5 + y) has two terms.

rebeccaxhawaii (rebeccaxhawaii):

What is (5+y)•8=?

OpenStudy (lfghjyt):

idk what is y

rebeccaxhawaii (rebeccaxhawaii):

Example, (x+7)•4 = 4•x+7•4= 4x+38

OpenStudy (lfghjyt):

whaaaaaa

OpenStudy (lfghjyt):

I need this can I just have the answer on this one?

OpenStudy (lfghjyt):

it has two factors and the factor (5+y) has two terms?